OpenStudy (anonymous):

Find the inverse function: f(x) = 6 - 3/(x^(2)) , x > 0

OpenStudy (anonymous):

y=6-3/x^2 Interchange x and y

OpenStudy (anonymous):

What am I doing wrong here? Here is my work so far: X = 6 - 3 / y^(2) x-6 = -3/y^(2) -1/3x +2 = y^2) (1/3x-2)^(1/2) = y

OpenStudy (anonymous):

can u?

OpenStudy (anonymous):

In step 3, I multiplied both sides by -1/3

hartnn (hartnn):

u multiplied by -1/3 but i took y^2 in numerator ....

OpenStudy (anonymous):

x-6 = -3/y^(2) y^2=(-3)/(x-6) y^2=-3/(x-6)

OpenStudy (anonymous):

Got your mistake?

OpenStudy (anonymous):

1 sec, let me look at what you wrote

OpenStudy (anonymous):

I dont get how you went from step 1 to 2

OpenStudy (anonymous):

x-6 = -3/y^(2) (x-6)*(y^2)=-3

OpenStudy (anonymous):

Do u understand it

OpenStudy (anonymous):

Ah yes.

OpenStudy (anonymous):

So now, you get y^(2) = -3 / x-6 y = square root(-3/x-6)

OpenStudy (anonymous):

y=+- square root ((-3)/(x-6))

OpenStudy (anonymous):

Just typed it on the website that I'm doing the homework on. The answer is correct :), tyvm

OpenStudy (anonymous):

Since x>0 for f(x), negative sign is rejected

OpenStudy (anonymous):

So it becomes y = square root(-3/x-6)

OpenStudy (anonymous):

yeah :). Thank you :D